### v1

# Tensor Decomposition Networks for Fast Machine Learning Interatomic Potential Computations

Abstract
$\rm{SO}(3)$-equivariant networks are the dominant models for machine learning interatomic potentials (MLIPs). The key operation of such networks is the Clebsch-Gordan (CG) tensor product, which is computationally expensive. To accelerate the computation, we develop tensor decomposition networks (TDNs) as a class of approximately equivariant networks whose CG tensor products are replaced by low-rank tensor decompositions, such as the CANDECOMP/PARAFAC (CP) decomposition. With the CP decomposition, we prove (i) a uniform bound on the induced error of $\rm{SO}(3)$-equivariance, and (ii) the universality of approximating any equivariant bilinear map. To further reduce the number of parameters, we propose path-weight sharing that ties all multiplicity-space weights across the $\mathcal{O}(L^{3})$ CG paths into a single path without compromising equivariance, where $L$ is the maximum angular degree. The resulting layer acts as a plug-and-play replacement for tensor products in existing networks, and the computational complexity of tensor products is reduced from $\mathcal{O}(L^{6})$ to $\mathcal{O}(L^{4})$. We evaluate TDNs on PubChemQCR, a newly curated molecular relaxation dataset containing 105 million DFT-calculated snapshots. We also use existing datasets, including OC20, and OC22. Results show that TDNs achieve competitive performance with dramatic speedup in computations.
2 Introduction
Symmetry is a fundamental aspect of molecular and material systems (zhang2023artificial, ), making it a crucial consideration in developing machine learning interatomic potentials (MLIPs). Equivariant graph neural networks have emerged as dominant frameworks in this domain. Usually, equivariant models incorporate directional features and spherical harmonics to maintain equivariance under rotation symmetry (gasteiger2019directional, ; satorras2021n, ; schutt2021equivariant, ). Among these, tensor product (TP) operations play a central role in fusing equivariant features, providing a powerful mechanism for building expressive models that adhere to $\rm{SO}(3)$-equivariance (thomas2018tensor, ; liao2022equiformer, ; batatia2022mace, ).
However, the computational cost of tensor product operations grows rapidly with the maximum angular degree $L$, reaching $\mathcal{O}(L^{6})$ in conventional implementations. Recent efforts to mitigate this cost have focused on accelerating the TP operation or applying frame averaging (FA) to enforce equivariance (duval2023faenet, ; dym2024equivariant, ; lin2024equivariance, ). While frame averaging is architecture-agnostic, it suffers from discontinuity issues (dym2024equivariant, ). On the other hand, TP acceleration techniques, such as SO(2)-based convolutions and fast spherical Fourier transformations, reduce complexity but still involve substantial parameter costs and computational overhead. Thus, there remains a gap in developing a method that simultaneously reduces computational complexity and parameter count while maintaining high accuracy and equivariance.
In this work, we propose Tensor Decomposition Networks (TDNs), a new class of approximately equivariant networks that replace the standard CG tensor product with low-rank tensor decompositions based on CANDECOMP/PARAFAC (CP) decomposition. TDNs introduce a CP decomposition that provides error bounds on equivariance and universality, ensuring consistency under $\rm{SO}(3)$ transformations. Additionally, a path-weight sharing mechanism consolidates multiplicity-space weights across CG paths, significantly reducing the parameter count from $\mathcal{O}(c^{2}L^{3})$ to $\mathcal{O}(c^{2})$ with $c$ the multiplicity. The resulting layer maintains the expressive power of conventional TP while lowering computational complexity from $\mathcal{O}(L^{6})$ to $\mathcal{O}(L^{4})$. We validate TDNs on a newly curated relaxation dataset with 105 million DFT-calculated molecular snapshots, along with the established OC20 and OC22 datasets, demonstrating competitive accuracy with substantial computational speedup.
3 Preliminaries and Related Work
Clebsch–Gordan (CG) tensor product is a fundamental operation widely used as the backbone for $\rm{SO}(3)$-equivariant neural network, enabling the fusion of feature fields at different angular degrees. We discuss the formal definition of CG tensor product with the maximum angular degree $L$. Consider two feature fields ${\bm{x}}=\bigoplus_{\ell_{1}=0}^{L}{\bm{x}}^{(\ell_{1})}$ and ${\bm{y}}=\bigoplus_{\ell_{2}=0}^{L}{\bm{y}}^{(\ell_{2})}$, where ${\bm{x}}^{(\ell_{1})},{\bm{y}}^{(\ell_{2})}$ are type-$\ell_{1}$ and type-$\ell_{2}$ irreducible representations (irreps) of $\rm{SO}(3)$. The CG tensor product fuses a pair of irreps ${\bm{x}}^{(\ell_{1})}$ and ${\bm{y}}^{(\ell_{2})}$ into every admissible output type $|\ell_{1}-\ell_{2}|\leq\ell_{3}\leq\ell_{1}+\ell_{2}$ via the CG coefficients $C^{\ell_{3},m_{3}}_{\ell_{1},m_{1},\ell_{2},m_{2}}$, defined as:
$({\bm{x}}^{(\ell_{1})}\!\otimes_{\!CG}\!{\bm{y}}^{(\ell_{2})})^{(\ell_{3})}_{m_{3}}=\sum_{m_{1}=-\ell_{1}}^{\ell_{1}}\sum_{m_{2}=-\ell_{2}}^{\ell_{2}}C^{\ell_{3},m_{3}}_{\ell_{1},m_{1},\ell_{2},m_{2}}\;{\bm{x}}^{(\ell_{1})}_{m_{1}}\,{\bm{y}}^{(\ell_{2})}_{m_{2}},\qquad-{\ell_{3}}\leq m_{3}\leq\ell_{3}.$
A triple $(\ell_{1},\ell_{2},\ell_{3})$ satisfying the CG selection rule $|\ell_{1}-\ell_{2}|\leq\ell_{3}\leq\ell_{1}+\ell_{2}$ is referred to as a path, and each path constitutes an independent $\rm{SO}(3)$-equivariant mapping. Collecting the contributions from all admissible paths, the complete CG tensor product is expressed as:
${\bm{x}}\otimes_{\!CG}{\bm{y}}=\bigoplus_{\ell_{1}=0}^{L}\;\bigoplus_{\ell_{2}=0}^{L}\;\bigoplus_{\ell_{3}=|\ell_{1}-\ell_{2}|}^{\ell_{1}+\ell_{2}\leq L}({\bm{x}}^{(\ell_{1})}\!\otimes_{\!CG}\!{\bm{y}}^{(\ell_{2})})^{(\ell_{3})}.$
However, the computational complexity of the CG tensor product scales as $\mathcal{O}(L^{6})$ as it involves $\mathcal{O}(L^{3})$ distinct paths and $\mathcal{O}(L^{3})$ operations per path. This significant cost poses a major bottleneck in practical implementations, especially when dealing with higher angular degrees. Computing the CG tensor product is very expensive due to the $\mathcal{O}(L^{6})$ computational complexity. To mitigate this computational challenge, inspired by tensor decomposition (kolda2009tensor, ) to low-rank tensors, we propose an efficient approximation based on tensor decomposition techniques, specifically the CP decomposition, to reduce the complexity while preserving approximate equivariance. The detailed formulation and implementation of the CP decomposition are presented in the following section.
Invariant and Equivariant Models. Symmetry has been a widely discussed constraint when developing machine learning methods for predicting chemical properties of molecules. Invariant and equivariant graph models have been widely applied in these cases. Invariant models (schutt2018schnet, ; gasteiger2019directional, ; gasteiger2021gemnet, ; liu2022spherical, ; wang2022comenet, ) aim to consider the rotation invariant features such as pairwise distance as the input, and take use of these to predict final properties. Equivariant models (satorras2021n, ; schutt2021equivariant, ; deng2021vector, ; jing2021learning, ; tholke2022equivariant, ) further incorporate equivariant features such as pairwise directions and spherical harmonics into the model. These models are built with equivariant blocks to ensure that output features rotate consistently with any rotation applied to the input features, thereby maintaining equivariant symmetry.
Tensor Product Acceleration. Among these equivariant networks, tensor product (thomas2018tensor, ; fuchs2020se, ; liao2022equiformer, ; batzner20223, ; batatia2022mace, ; unke2021se, ; yu2023efficient, ) is one of the most important components that fuse two equivariant features into one. It provides a powerful and expressive way (dym2020universality, ) to build equivariant networks, while the computational cost of TP is usually considerable. Therefore, there are several directions to accelerate the equivariant networks. First direction is to accelerate TP. eSCN (liao2024equiformerv, ; passaro2023reducing, ) proposes to reduce the SO(3) convolution into SO(2) for TP when one of the inputs of TP is spherical harmonics. Gaunt tensor product (luoenabling, ) makes use of fast spherical Fourier transformation to perform the TP. The other direction is to apply frame averaging (FA) (duval2023faenet, ; dym2024equivariant, ; lin2024equivariance, ), which uses group elements from an equivariant set-value function called frame to transform the input data and subsequently the model’s output, enabling any models to obtain the desired symmetries. Although it is flexible and has no requirement for model architectures, it faces an unsolvable discontinuity problem (dym2024equivariant, ).
4 Tensor Decomposition Networks
This section presents the techniques employed in the proposed Tensor Decomposition Network (TDN). In Section 4.1, we introduce a CP-decomposition-based approximation for the tensor product, and in Section 4.2, we detail a path-wise weight-sharing scheme. Together, these strategies effectively reduce both computational cost and parameter count. In Section 4.3, we analyze the computational complexity of the approximate tensor product, and in Section 4.4, we discuss its error bound and universality.
4.1 Tensor Product and Its Approximation
The tensor product is a fundamental operation in equivariant neural networks, enabling the coupling of features across multiple vector spaces. However, direct implementation of the tensor product incurs significant computational cost. To mitigate this, we introduce a low-rank approximation using the CANDECOMP/PARAFAC (CP) decomposition to reduce the time complexity.
Tensor Product.
Before developing our tensor-product approximation we recall the canonical definition of the tensor product in the simplest non-trivial two-order case. The multi-order counterpart and its CP decomposition are introduced in Appendix A. In practice, higher-rank tensors are stored as flattened vectors via a fixed index ordering; this reshaping preserves the vector-space operations. Without loss of generality, we present the following tensor product definition. Let $V_{1}=\mathbb{R}^{d_{1}}$, $V_{2}=\mathbb{R}^{d_{2}}$, and $V_{3}=\mathbb{R}^{d_{3}}$ be finite-dimensional real vector spaces with fixed ordered bases $\{{\bm{e}}_{i}\}_{i=1}^{d_{1}},\{{\bm{f}}_{j}\}_{j=1}^{d_{2}},\{{\bm{g}}_{k}\}_{k=1}^{d_{3}}$. The tensor product $V_{1}\otimes V_{2}$ is the space that corepresents bilinear maps: for every bilinear $m\colon V_{1}\times V_{2}\to V_{3}$ there exists a unique linear map $\widetilde{m}\colon V_{1}\otimes V_{2}\;\longrightarrow\;V_{3}$ such that $m({\bm{x}},{\bm{y}})=\widetilde{m}({\bm{x}}\otimes{\bm{y}})$. With respect to the chosen bases, $\widetilde{m}$ is encoded by a three–way tensor
${\bm{M}}=\bigl{(}{\bm{M}}_{kij}\bigr{)}\;\in\;V_{3}\otimes V_{1}^{\!*}\otimes V_{2}^{\!*}\;\cong\;\mathbb{R}^{d_{3}\times d_{1}\times d_{2}},$
and $m({\bm{e}}_{i},{\bm{f}}_{j})=\sum_{k=1}^{d_{3}}{\bm{M}}_{kij}\,{\bm{g}}_{k}$. For arbitrary ${\bm{x}}=\sum_{i=1}^{d_{1}}x_{i}{\bm{e}}_{i}$ and ${\bm{y}}=\sum_{j=1}^{d_{2}}y_{j}{\bm{f}}_{j}$ we have
$m({\bm{x}},{\bm{y}})=\sum_{k=1}^{d_{3}}\sum_{i=1}^{d_{1}}\sum_{j=1}^{d_{2}}{\bm{M}}_{kij}\,x_{i}\,y_{j}\,{\bm{g}}_{k}={\bm{M}}\bigl{(}{\bm{x}}\otimes{\bm{y}}\bigr{)}.$ (1)
CP Decomposition for Tensor Product Approximation.
To reduce time complexity and parameter count of tensor product, we approximate the tensor ${\bm{M}}$ via CP decomposition (kolda2009tensor, ) of rank $R$. The CP decomposition writes the tensor product as sum of $R$ rank-1 tensors by decomposing the three-way tensor ${\bm{M}}$ such that
${\bm{M}}_{kij}\approx\sum_{r=1}^{R}{\bm{A}}_{kr}{\bm{B}}_{ir}{\bm{C}}_{jr},$ (2)
where ${\bm{A}}\in\mathbb{R}^{d_{3}\times R}$, ${\bm{B}}\in\mathbb{R}^{d_{1}\times R}$, and ${\bm{C}}\in\mathbb{R}^{d_{2}\times R}$ are factor matrices capturing the modes of the tensor. Substituting Eq. 2 into Eq. 1 gives
$m({\bm{x}},{\bm{y}})\approx\sum_{k=1}^{d_{3}}\sum_{i=1}^{d_{1}}\sum_{j=1}^{d_{2}}\left(\sum_{r=1}^{R}{\bm{A}}_{kr}{\bm{B}}_{ir}{\bm{C}}_{jr}\right)x_{i}y_{j}{\bm{g}}_{k}.$
Then we rearrange the summation to obtain
$\sum_{r=1}^{R}\left(\sum_{i=1}^{d_{1}}{\bm{B}}_{ir}x_{i}\right)\left(\sum_{j=1}^{d_{2}}{\bm{C}}_{jr}y_{j}\right)\left(\sum_{k=1}^{d_{3}}{\bm{A}}_{kr}{\bm{g}}_{k}\right).$
For clarity, the above approximation for the order-two tensor product $m({\bm{x}},{\bm{y}})$ can be expressed in matrix form as:
$m({\bm{x}},{\bm{y}})\approx{\bm{A}}\bigl{(}{\bm{B}}^{\top}{\bm{x}}\odot{\bm{C}}^{\top}{\bm{y}}\bigr{)},$ (3)
where $\odot$ denotes the Hadamard product. By the universal property that every bilinear map $m\colon V_{1}\times V_{2}\to V_{3}$ factors uniquely as a linear map $\,\widetilde{m}\colon V_{1}\otimes V_{2}\to V_{3}$, the tensor product covers important instances used in equivariant learning, notably the CG tensor product (thomas2018tensor, ) and the Gaunt tensor product (luoenabling, ). Therefore, the tensor product approximation can be employed for those specific cases. In this paper, we primarily discuss CG tensor product approximation.
Figure 1: Illustration of approximating the CG tensor product using CP decomposition. The input features $\bm{x}$ and $\bm{y}$ both consist of irreps with $\ell=1$, and the CG tensor product produces output feature $\bm{z}$ containing irreps with $\ell=0,1,2$.
Figure 2: Approximation error for different rank values $R$ across various maximum angular degrees $L$.
CP Decomposition for CG Tensor Product.
Next, we introduce the idea to make use of CP decomposition to accelerate the CG tensor product calculations. Specifically, the CG coefficient tensor is a three-way tensor concatenating CG coefficients ${\bm{C}}_{\ell_{1},\ell_{2}}^{\ell_{3}}$ of all admissible path $(\ell_{1},\ell_{2},\ell_{3})$ such that
${\bm{M}}=\bigoplus_{\ell_{1}=0}^{L}\bigoplus_{\ell_{2}=0}^{L}\bigoplus_{\ell_{3}=|\ell_{1}-\ell_{2}|}^{\ell_{1}+\ell_{2}\leq L}{\bm{C}}_{\ell_{1},\ell_{2}}^{\ell_{3}}.$
The key idea of CP decomposition is to break down the CG coefficient tensor ${\bm{M}}$ into low-rank matrices. We demonstrate a simple example of CP decomposition for TP with $\ell_{1},\ell_{2}=1$ and $\ell_{3}=(0,1,2)$, as shown in Fig. 1. Eq. 3 allows for the simple batch-wise use of the Hadamard product with a global hyperparameter $R$ instead of computing tensor product per path. The higher $R$, the more accurate the tensor product result is. For acceleration, we do not need full rank $R$ and we can use a small $R$ while keeping a low approximation error. We show the rank-error curve in Fig. 2.
4.2 Path-Weight Sharing Tensor Product
In equivariant neural networks, the CG tensor product couples features that transform under irreducible representations (irreps) of $\mathrm{SO}(3)$. For a maximal degree $L$, a fully connected CG tensor product introduces $\mathcal{O}(L^{3})$ paths, each associated with a distinct weight matrix. This leads to a substantial parameter count, which can hinder model efficiency. To mitigate this, we propose a path-weight sharing mechanism to reduce the parameter count while retaining equivariance.
Concatenating Irreps into a Single Channel.
We first unify the multiplicities across all degrees $\ell$ by setting a common multiplicity $c$. Letting ${\bm{x}}^{(\ell)}\in\mathbb{R}^{c\times(2\ell+1)}$ denote the multiplicity-$c$ irrep of degree $\ell$, we concatenate the multiplicity axes into a single channel:
$\tilde{\bm{x}}=\operatorname{concat}_{\ell=0}^{L}{\bm{x}}^{(\ell)}\in\mathbb{R}^{c\times(L+1)^{2}}.$
All irreps now share a common multiplicity index, so the linear projection and CG contraction over irreps features are executed as batched matrix operations whose operands reside contiguously in memory; this yields coalesced memory access on GPUs and improved cache locality on CPUs.
Path-Weight Sharing Tensor Product.
We further reduce the parameter count by applying a path-weight sharing scheme. Let ${\bm{W}}_{\ell_{1},\ell_{2}}^{\ell}\in\mathbb{R}^{c\times c}$ denote the multiplicity-space weight matrix associated with the path $(\ell_{1},\ell_{2},\ell_{3})$, where $0\leq\ell_{1},\ell_{2},\ell_{3}\leq L$ and $c$ is the channel dimension of each irrep block. Instead of assigning a unique matrix for each path, we set all such matrices equal to a single learnable parameter ${\bm{W}}\in\mathbb{R}^{c\times c}$, i.e. ${\bm{W}}_{\ell_{1},\ell_{2}}^{\ell}\equiv{\bm{W}}$ for every admissible path $(\ell_{1},\ell_{2},\ell_{3})$. This collapses the $\mathcal{O}(L^{3})$ distinct weight tensors of the naïve implementation into one, reducing the parameter count from $\mathcal{O}\bigl{(}c^{2}L^{3}\bigr{)}$ to $\mathcal{O}\bigl{(}c^{2}\bigr{)}$. Because the sharing operates exclusively on multiplicity indices, the irrep content of each block and hence full $\mathrm{SO}(3)$ equivariance is retained. The resulting layer therefore generalizes the classic CG tensor product while offering an orders-of-magnitude reduction in parameters.
4.3 Complexity Analysis of Approximate Tensor Product
In this section, we analyze the computational complexity of the proposed approximate tensor product using CP decomposition. We first investigate the rank of the tensor product. Let $\operatorname{rank}_{\mathrm{CP}}({\bm{M}})$ denote the CP rank of the three-way tensor ${\bm{M}}\in\mathbb{R}^{d_{3}\times d_{1}\times d_{2}}$, representing the minimal rank $R$ such that there is an equality for Eq. 2, i.e.
$\operatorname{rank}_{\mathrm{CP}}({\bm{M}})=\min\Bigl{\{}R\in\mathbb{N}^{+}\;\Bigm{|}\;{\bm{M}}_{kij}=\sum_{r=1}^{R}{\bm{A}}_{kr}{\bm{B}}_{ir}{\bm{C}}_{jr}\Bigr{\}}.$
Determining $\operatorname{rank}_{\mathrm{CP}}({\bm{M}})$ exactly is NP-hard (kolda2009tensor, ). In practice one specifies a rank $R$ as a hyper-parameter and optimizes the factor matrices ${\bm{A}}\in\mathbb{R}^{d_{3}\times R}$, ${\bm{B}}\in\mathbb{R}^{d_{1}\times R}$, ${\bm{C}}\in\mathbb{R}^{d_{2}\times R}$ to minimize a prescribed loss. A generic upper bound $\operatorname{rank}_{\mathrm{CP}}({\bm{M}})\;\leq\;\min\{d_{1}d_{2},\;d_{1}d_{3},\;d_{2}d_{3}\}$ limits the choices of $R$. Consequently, algorithmic pipelines treat $R$ as an external choice, balancing approximation accuracy against computational cost.
By using CP decomposition, the computational cost for evaluating the approximate tensor product in Eq. 3 reduces to $\mathcal{O}\bigl{(}R(d_{1}+d_{2}+d_{3})\bigr{)}$, a significant reduction compared to the full tensor contraction cost of $\mathcal{O}(d_{1}d_{2}d_{3})$ using Eq. 1. Similarly, the parameter budget decreases from $\mathcal{O}(d_{1}d_{2}d_{3})$ to $\mathcal{O}(R(d_{1}+d_{2}+d_{3}))$, which provides substantial savings when $R$ is small.
For the approximation of the CG tensor product, where $d_{1},d_{2},d_{3}\propto\mathcal{O}(L^{2})$, the computational complexity further reduces to $\mathcal{O}(RL^{2})$. In our experiments, we select $R=7L^{2}$. The error curve for the CP decomposition with varying $R$ is discussed in Section 5.1.
4.4 Error Bound and Universality Analysis
The error bound and universality analysis provide theoretical guarantees for the CP decomposition of the tensor product in the proposed Tensor Decomposition Network (TDN). This section establishes the error bound for both the approximation and equivariance, demonstrating how the approximation error depends on the spectral tail of the tensor’s singular values. Additionally, we formalizes the universality property of the CP decomposition, showing that as the rank $R$ increases, the CP-decomposition-based tensor product can accurately approximate any $\mathrm{SO}(3)$-equivariant bilinear map, thereby preserving the expressive power of the tensor product while reducing computational complexity.
Error Bound of Approximate Tensor Product.
Given a rank $R\leq\operatorname{rank}_{\mathrm{CP}}({\bm{M}})$, one seeks CP-decomposition-based approximation $\widehat{\bm{M}}$ by minimizing $\lVert{\bm{M}}-\widehat{\bm{M}}\rVert$ in a chosen norm, typically the Frobenius norm. Although the non-convex optimization may admit spurious local minima, the optimization of CP decomposition possesses an essentially unique best rank-$R$ approximation under Kruskal’s condition (kolda2009tensor, ), and modern alternating least-squares or gradient methods converge to it under mild coherence assumptions (yang2023global, ). For error estimation in our setting, we specialize to the CG tensor product, where we let all irreps have the same maximum degree and all dimensions equal such that $d=d_{1}=d_{2}=d_{3}$. Given the singular values $\sigma_{k}^{(n)}$ of the mode-$n$ matricization ${\bm{M}}_{(n)}$ of ${\bm{M}}$ and truncating each ${\bm{M}}_{(n)}$ to rank $R_{T}=\lceil R^{1/3}\rceil$, a priori approximation error bound (de2000multilinear, ) gives
$\bigl{\lVert}{\bm{M}}-\widehat{{\bm{M}}}\bigr{\rVert}_{F}\;\leq\;\Bigl{(}\sum_{n=1}^{3}\sum_{k>R_{T}}\sigma_{k}^{(n)2}\Bigr{)}^{1/2}.$
To quantify the loss of $\mathrm{SO}(3)$-equivariance incurred by CP decomposition, let ${\bm{R}}\in\mathrm{SO}(3)$ with representation the Wigner D-matrix $D({\bm{R}})$, and the $\rm{SO}(3)$-equivariance error is estimated by
$\varepsilon({\bm{R}},{\bm{x}},{\bm{y}})=\lVert\widehat{{\bm{M}}}(D({\bm{R}}){\bm{x}}\otimes D({\bm{R}}){\bm{y}})-D({\bm{R}})\widehat{{\bm{M}}}({\bm{x}}\otimes{\bm{y}})\rVert.$
The following result bounds this error uniformly over $\rm{SO}(3)$ with the proof in Section B.1. {theorem}Equivariance Error Bound of CP Decompositionequivariance-error Let CG tensor ${\bm{M}}\in\mathbb{R}^{d\times d\times d}$ and $\widehat{{\bm{M}}}$ be the rank-$R$ CP-decomposition-based approximation obtained by Frobenius minimization. For any rotation ${\bm{R}}\in SO(3)$ and any bounded irreps ${\bm{x}},{\bm{y}}\in\mathbb{R}^{d},\lVert{\bm{x}}\rVert,\lVert{\bm{y}}\rVert\leq C$, we have
[…]
where $R_{T}=\lceil R^{1/3}\rceil$ and $\sigma_{k}^{(n)}$ is the $k$-th singular value of mode-$n$ matricization of ${\bm{M}}$. Empirical estimates of both the approximation and equivariance errors are provided in Section 5.1.
Universality of Approximate Tensor Product.
Because the tensor product is universal for bilinear maps, any $\rm{SO}(3)$-equivariant bilinear operator can be expressed as a composition of a tensor product followed by a suitable ${\bm{M}}$ onto an equivariant subspace. Consequently, the CP-decomposition-based approximation developed above inherits this universality: for every equivariant tensor product there exists a rank-$R$ approximation that converges to it as $R\to\operatorname{rank}_{\mathrm{CP}}({\bm{M}})$. The following theorem formalizes the expressivity of our approximation scheme with the proof in Section B.2.
{theorem}
Universality of CP Decompositionuniversality Consider $\rm{SO}(3)$-irreps ${\bm{x}}\in V_{1}=\mathbb{R}^{d}$, ${\bm{y}}\in V_{2}=\mathbb{R}^{d_{2}}$ and $V_{3}=\mathbb{R}^{d_{3}}$. Fo any $\rm{SO}(3)$-equivariant bilinear map $\Phi$, there exist ${\bm{B}}\in\mathbb{R}^{d_{1}\times R}$, ${\bm{C}}\in\mathbb{R}^{d_{2}\times R}$, ${\bm{A}}\in\mathbb{R}^{R\times d_{3}}$ such that $\Phi$ can be written as
$\Phi({\bm{x}},{\bm{y}})={\bm{A}}({\bm{B}}^{\top}{\bm{x}}\odot{\bm{C}}^{\top}{\bm{y}})\in V_{3},$
if $R$ is sufficient large.
5 Experiments
The effectiveness of our method is assessed on two benchmarks: the PubChemQCR dataset, which contains millions of molecular-dynamics snapshots, and the established Open Catalyst Project (OCP) datasets. Section 5.1 describes both benchmarks and model configurations. Section 5.2 then analyses the computational efficiency of our approach as the maximum angular degree in the tensor product is varied. Section 5.3 compares our model with several baselines on PubChemQCR and its subset PubChemQCR-S, while Section 5.4 and Section 5.5 report results of our model compared to several tensor-product baselines on OC20 and OC22 from OCP, respectively.
5.1 Experimental Setting
Table 1: Comparison of model performance on energy and force predictions for the PubChemQCR and the PubChemQCR-S dataset. Our model is trained to compare against several baseline methods on the PubChemQCR-S dataset, including SchNet (schutt2018schnet, ), PaiNN (schutt2021equivariant, ), MACE (batatia2022mace, ), PACE (xu2024equivariant, ), FAENet (duval2023faenet, ), NequIP (batzner20223, ), SevenNet (park2024scalable, ), Allegro (musaelian2023learning, ), and Equiformer (liao2022equiformer, ), and. On the full PubChemQCR dataset, we compare performance with SchNet (schutt2018schnet, ) and PaiNN (schutt2021equivariant, ).
Validation Test
Dataset Model Energy MAE Force MAE Energy MAE Force MAE
(meV) $\downarrow$ (meV/Å) $\downarrow$ (meV) $\downarrow$ (meV/Å) $\downarrow$
Small Subset SchNet $5.30$ $56.55$ $5.55$ $56.22$
PaiNN $5.13$ $46.34$ $5.33$ $46.92$
NequIP $7.37$ $54.78$ $8.27$ $55.59$
SevenNet $8.77$ $47.63$ $10.21$ $48.05$
Allegro $10.86$ $60.71$ $10.80$ $61.44$
FAENet $7.28$ $60.24$ $8.70$ $60.51$
MACE $7.54$ $51.46$ $7.47$ $45.70$
PACE $6.24$ $50.54$ $6.53$ $51.43$
Equiformer 4.69 34.67 5.38 35.11
TDN 4.46 26.94 5.01 26.43
Full SchNet 7.14 65.22 7.71 67.38
PaiNN 3.62 38.30 3.49 39.28
TDN 1.65 19.46 1.50 20.44
Datasets.
We use a newly curated dataset PubChemQCR (fu2025benchmarkquantumchemistryrelaxations, ), comprising high-fidelity molecular trajectories derived from the PubChemQC database (nakata2017pubchemqc, ). This dataset encompasses a diverse range of molecular systems, capturing potential energy surfaces and force information critical for understanding molecular interactions. The full dataset consists of 3,471,000 trajectories and 105,494,671 DFT-calculated molecular snapshots, with each snapshot containing molecular structure, total energy, and forces. For training efficiency, we use the smaller subset, PubChemQCR-S, comprising 40,979 trajectories and 1,504,431 molecular snapshots for model benchmarking. The subset is split into training, validation, and testing sets using a 60%-20%-20% ratio, while the full dataset employs an 80%-10%-10% split to assess generalizability. To prevent data leakage, each trajectory is assigned to only one split in both PubChemQCR and PubChemQCR-S.
The second dataset used in this study is the Open Catalyst Project (OCP) (zitnick2020introduction, ), which publishes extensive open datasets of DFT relaxations for adsorbate–catalyst surfaces and hosts public leaderboard challenges. The flagship releases, OC20 (chanussot2021open, ) and OC22 (tran2023open, ), encompass thousands of chemical compositions, crystal facets, and adsorbates, serving as comprehensive benchmarks for surrogate models aiming to replace computationally intensive quantum-mechanical calculations. Each release defines multiple tasks, including Initial-Structure-to-Relaxed-Energy (IS2RE) and Structure-to-Energy-and-Force (S2EF), which require accurate predictions of total energies, per-atom forces, and relaxed geometries. We conduct experiments on OC20 IS2RE and S2EF, as well as OC22 IS2RE, using the data splits and configurations specified in the official OCP repository.
(a) Approximation error vs. $L$
(b) $\rm{SO}(3)$-equivariance error vs. $L$
(c) Runtime vs. $L$ (log–scale)
Figure 3: Scaling behaviours of the CP-decomposition-based tensor product under different maximum angular degree scheduler: (a) approximation error, (b) $\rm{SO}(3)$-equivariance error, and (c) runtime.
CP Rank Selection.
We approximate the fully connected CG tensor product using the rank-$R$ CP decomposition across various configurations of the maximum angular degree $L$. Four rank schedules are evaluated: $R=(L+1)^{4}$, $R=16\log^{2}(L+1)$, $R=7L$, and $R=7L^{2}$. For each configuration we measure (i) Approximation Error: The relative CP tensor product error, calculated as $\lVert{\bm{M}}({\bm{x}}\otimes{\bm{y}})-\widehat{{\bm{M}}}({\bm{x}}\otimes{\bm{y}})\rVert_{F}/\lVert{\bm{M}}({\bm{x}}\otimes{\bm{y}})\rVert_{F}$; (ii) Equivariance Error: The expected $\mathrm{SO}(3)$-equivariance error, defined as $\mathbb{E}_{{\bm{R}},{\bm{x}},{\bm{y}}}\bigl{[}\varepsilon({\bm{R}},{\bm{x}},{\bm{y}})\bigr{]}$, where the expectation is averaged over $1000$ random rotations ${\bm{R}}$ and random vectors ${\bm{x}},{\bm{y}}$; and (iii) Execution Time: The runtime for the CP-decomposition-based tensor product under each rank schedule. Our results demonstrate that the logarithmic schedule $R=7L^{2}$ consistently achieves the lowest approximation and equivariance errors across all $L$, while requiring significantly less computation time compared to the CG tensor product baseline. Thus, this schedule provides the best balance between accuracy and computational cost and is adopted for all subsequent experiments.
Model Design.
Building on the capabilities of graph transformers, we develop the Tensor-Decomposition Network (TDN) by modifying the publicly available Equiformer architecture (liao2022equiformer, ). Specifically, we replace every channel-wise linear projection, normalization layer, and activation in Equiformer with batched matrix operations that act simultaneously on the multiplicity dimension of each irrep block, eliminating the need for costly slicing and reshaping between tensor and vector representations. Additionally, the core CG tensor products within the self-attention mechanism are substituted with our rank-$R$ CP-decomposition-based tensor product from Section 4.1, integrated with the path-weight-sharing scheme from Section 4.2. This design preserves the expressive power of Equiformer while significantly improving memory and computational efficiency. The model and training configurations for TDN are described in Appendix C.
Baseline Implementations.
For the PubChemQCR benchmark, we reimplement each baseline model based on its official repository. Hyperparameters are adopted from the best configurations reported in the original papers or, if unspecified, are tuned using the PubChemQCR-S dataset. All model and training configurations of baseline models are described in Appendix C.
5.2 Efficiency of CP-Decomposition-Based Tensor Product
Table 2: Throughput and parameter count comparison between TDN and Equiformer across different values of maximum degree $L$.
Model Throughput (samples/sec) Parameter
Equiformer (L=1) 470.9 8.86M
TDN (L=1) 583.1 4.42M
Equiformer (L=2) 122.5 18.19M
TDN (L=2) 199.7 4.42M
Equiformer (L=3) 48.4 33.43M
TDN (L=3) 142.9 4.42M
To evaluate the speed-up achieved by our CP-decomposition-based tensor product in Section 4.1, we benchmark its wall-clock runtime against the fully connected CG tensor product implementation in e3nn. As shown in Fig. 3(c), the proposed approximation accelerates by factors of 4.0×,4.8×,15.0×,26.7×,47.6×,4.0$\times$,4.8$\times$,15.0$\times$,26.7$\times$,47.6$\times$,4.0 × , 4.8 × , 15.0 × , 26.7 × , 47.6 × , and 83.6×83.6$\times$83.6 × for maximum degrees $L=1,2,3,4,5,$ and $6$, respectively.
Since TDN is derived from Equiformer by replacing each CG block with its CP decomposition and incorporating path-weight sharing, we further evaluate end-to-end throughput and parameter count for both networks under identical irrep configurations across varying $L$. All experiments are conducted on a single NVIDIA A100-80GB GPU with batch size 128 and identical input irreps. The results in Table 2 show that TDN processes 1.23×1.23$\times$1.23 × to 2.95×2.95$\times$2.95 × more structures per second and uses $50\%$$87\%$ fewer parameters than Equiformer. As the maximum degree $L$ increases, TDN achieves higher throughput and requires fewer parameters.
5.3 Results on PubChemQCR
We evaluate our method on both PubChemQCR-S and the full PubChemQCR dataset. For the small split we compare against nine state-of-the-art models: SchNet (schutt2018schnet, ), PaiNN (schutt2021equivariant, ), MACE (batatia2022mace, ), PACE (xu2024equivariant, ), FAENet (duval2023faenet, ), NequIP (batzner20223, ), SevenNet (park2024scalable, ), Allegro (musaelian2023learning, ), and Equiformer (liao2022equiformer, ). On the full dataset we benchmark against SchNet and PaiNN, the only baselines that scale to its size within our hardware budget. Performance is reported as mean absolute error (MAE) for energies and root-mean-square error (RMSE) for forces. The results are shown in Table 1, the proposed TDN model achieves the lowest energy and force prediction errors across PubChemQCR-S and the full PubChemQCR dataset, significantly outperforming all baseline methods. Moreover, the performance further improves as the dataset size increases.
5.4 Results on OC20
Table summarizes our performance on the two principal OC20 validation tasks: IS2RE-Direct, which predicts the relaxed adsorption energy directly from the initial geometry (no noisy-node auxiliary loss), and S2EF on the 2M subset, which jointly regresses per-atom forces and total energies.
Table 3: Comparison of model performance on energy and force predictions for OC20 S2EF 2M validation set. Our model is trained to compare against several baseline methods, including SchNet (schutt2018schnet, ), DimeNet++ (gasteiger2020fast, ), SpinConv (shuaibi2021rotation, ), GemNet-OC (gasteiger2022gemnet, ), eSCN (passaro2023reducing, ), and EquiformerV2 (liao2024equiformerv, ).
Model Energy MAE Force MAE
(meV) (meV/Å)
SchNet 1400 78.3
DimeNet++ 805 65.7
SpinConv 406 36.2
GemNet-OC 286 25.7
eSCN 283 20.5
EquiformerV2 285 20.5
TDN 257 30.8
We benchmark against the strongest tensor-product architectures reported to date, including Equiformer (liao2022equiformer, ), Equiformer-v2 (liao2024equiformerv, ), and eSCN (passaro2023reducing, ). Metrics follow the official OC20 protocol: energy mean absolute error (MAE, eV) and energy within threshold (EwT, %) in IS2RE-Direct for four validation sub-splits: distribution adsorbates and catalysts (ID), out-of-distribution adsorbates (OOD-Ads), out-of-distribution catalysts (OOD-Cat), and out-of-distribution adsorbates and catalysts (OOD-Both), and energies MAE (meV) and forces MAE (meV/Å) in S2EF 2M validation set. The S2EF results are shown in Table 3. Our TDN model achieves superior performance in energy prediction while maintaining comparable accuracy in force prediction. The IS2RE-Direct results are presented in Table 4, demonstrating that our model achieves performance comparable to Equiformer while being more efficient, as detailed in Section 5.2. This highlights the effectiveness of our architecture in maintaining expressivity while significantly reducing computational costs.
Both the IS2RE-Direct and S2EF 2M TDN models are trained with the optimization scheme of Equiformer (liao2022equiformer, ): AdamW with learning-rate $2\times 10^{-4}$ and weight decay $10^{-3}$, a batch size of 32, and a cosine decay schedule with 2 warm-up epochs and 0.2 warm-up factor over 30 training epochs.
Table 4: Comparison of model performance on energy predictions for OC20 IS2RE-Direct validation set without noisy-node auxiliary loss. Our model is trained to compare against several baseline methods, including SchNet (schutt2018schnet, ), DimeNet++ (gasteiger2020fast, ), GemNet-dT (gasteiger2021gemnet, ), SphereNet (liu2022spherical, ), Equiformer (liao2022equiformer, ) and EquiformerV2 (liao2024equiformerv, ).
Energy MAE (eV) $\downarrow$ EwT (%) $\uparrow$
Model ID OOD Ads OOD Cat OOD Both Average ID OOD Ads OOD Cat OOD Both Average
SchNet 0.6465 0.7074 0.6475 0.6626 0.6660 2.96 2.22 3.03 2.38 2.65
DimeNet++ 0.5636 0.7127 0.5612 0.6492 0.6217 4.25 2.48 4.40 2.56 3.42
GemNet-dT 0.5561 0.7342 0.5659 0.6964 0.6382 4.51 2.24 4.37 2.38 3.38
SphereNet 0.5632 0.6682 0.5590 0.6190 0.6024 4.56 2.70 4.59 2.70 3.64
Equiformer 0.5088 0.6271 0.5051 0.5545 0.5489 4.88 2.93 4.92 2.98 3.93
EquiformerV2 0.5161 0.7041 0.5245 0.6365 0.5953 - - - - -
TDN 0.5085 0.6668 0.5104 0.5875 0.5683 5.21 2.54 5.04 2.98 3.94
5.5 Results on OC22
Table 5: Comparison of model performance on energy predictions for OC22 IS2RE-Direct testing set. We compare with several baseline methods, including SchNet (schutt2018schnet, ), DimeNet++ (gasteiger2020fast, ), PaiNN (schutt2021equivariant, ), GemNet-dT (gasteiger2021gemnet, ), and coGN (ruff2302connectivity, ).
Model MAE (ID) MAE (OOD) Average
(eV) (eV) (eV)
SchNet 2.00 4.85 3.42
DimeNet++ 1.96 3.52 2.74
PaiNN 1.72 3.68 2.70
GemNet-dT 1.68 3.08 2.38
coGN 1.62 2.81 2.21
TDN 1.49 2.92 2.20
Following the OC20 protocol, we assess our model on the OC22 IS2RE-Direct test set, in which the energy is predicted directly from the initial structure without the noisy-node auxiliary loss. Metrics include the average, in-distribution (ID) and out-of-distribution (OOD) mean absolute error (MAE, eV) obtained from the four standard OC22 splits with the same split protocol as mentioned in OC20. For training, the TDN model is trained with the same optimization scheme as for OC20 IS2RE-Direct but for 1000 epochs. As shown in Table 5, TDN achieves the lowest in-distribution and average MAE and second best performance on out-of-distribution MAE, validating the high effectiveness of TDN.
6 Summary
In this work, we present Tensor Decomposition Networks (TDNs), a novel framework designed to accelerate the computationally intensive CG tensor product in $\rm{SO}(3)$-equivariant networks through low-rank tensor decomposition. By leveraging CANDECOMP/PARAFAC (CP) decomposition and implementing path-weight sharing, TDNs effectively reduce both parameter count and computational complexity while preserving the expressive power of conventional tensor products. Extensive evaluations on a newly curated PubChemQCR dataset and commonly used OC20 and OC22 benchmarks demonstrate that TDNs achieve comparable predictive accuracy to state-of-the-art models while significantly reducing runtime. The proposed framework provides a plug-and-play alternative to conventional tensor products, making it a promising approach for large-scale molecular simulations and MLIP training.
Acknowledgments
SJ acknowledges support from ARPA-H under grant 1AY1AX000053, National Institutes of Health under grant U01AG070112, and National Science Foundation under grant IIS-2243850. XFQ acknowledges support from the Air Force Office of Scientific Research (AFOSR) under Grant FA9550-24-1-0207. We acknowledge the support of Lambda, Inc. and NVIDIA for providing the computational resources for this project.
Appendix A Multilinear Maps and Tensor Decomposition in Arbitrary Order
Tensor product in arbitrary order.
Let $N\geq 2$ and let $V_{i}=\mathbb{R}^{d_{i}}$ for $i=1,\dots,N+1$ be finite-dimensional real vector spaces equipped with fixed ordered bases $\{{\bm{e}}_{k}^{(i)}\}_{k=1}^{d_{i}}$. A multilinear map
$m\colon V_{1}\times\cdots\times V_{N}\longrightarrow V_{N+1}$
is uniquely represented by a linear map
$\widetilde{m}\colon V_{1}\otimes\cdots\otimes V_{N}\;\longrightarrow\;V_{N+1},\qquad m({\bm{x}}^{(1)},\dots,{\bm{x}}^{(N)})=\widetilde{m}\!\bigl{(}{\bm{x}}^{(1)}\otimes\cdots\otimes{\bm{x}}^{(N)}\bigr{)}.$
With respect to the chosen bases, $\widetilde{m}$ is encoded by an $(N\!+\!1)$-way tensor
${\bm{M}}\in V_{N+1}\otimes V_{1}^{\!*}\otimes\cdots\otimes V_{N}^{\!*}\cong\mathbb{R}^{d_{N+1}\times d_{1}\times\cdots\times d_{N}},$
defined through
$m\!\bigl{(}{\bm{x}}^{(1)},\dots,{\bm{x}}^{(N)}\bigr{)}\;=\;\sum_{i_{N+1}=1}^{d_{N+1}}\sum_{i_{1}=1}^{d_{1}}\cdots\sum_{i_{N}=1}^{d_{N}}{\bm{M}}_{i_{N+1}i_{1}\cdots i_{N}}{\bm{x}}^{(1)}_{i_{1}}\cdots{\bm{x}}^{(N)}_{i_{N}}{\bm{e}}^{(N+1)}_{i_{N+1}}.$ (4)
CP decomposition in arbitrary order.
To lower computational cost, we can approximate ${\bm{M}}$ by a rank-$R$ CP decomposition
${\bm{M}}_{i_{N+1}i_{1}\cdots i_{N}}\;\approx\;\sum_{r=1}^{R}{\bm{A}}_{i_{N+1}r}\,{\bm{B}}^{(1)}_{i_{1}r}\cdots{\bm{B}}^{(N)}_{i_{N}r},$
where ${\bm{A}}\in\mathbb{R}^{d_{N+1}\times R}$ and ${\bm{B}}^{(i)}\in\mathbb{R}^{d_{i}\times R}\;(i=1,\dots,N)$. Substituting into Eq. 4 and rearranging the summation the we obtain
$m\!\bigl{(}{\bm{x}}^{(1)},\dots,{\bm{x}}^{(N)}\bigr{)}\;\approx\;\sum_{r=1}^{R}\left(\sum_{i_{1}=1}^{d_{1}}{\bm{B}}_{i_{1}r}^{(1)}{\bm{x}}_{i_{1}}^{(1)}\right)\cdots\left(\sum_{i_{N}=1}^{d_{N}}{\bm{B}}_{i_{N}r}^{(N)}{\bm{x}}_{i_{N}}^{(N)}\right)\left(\sum_{i_{N+1}=1}^{d_{N+1}}{\bm{A}}_{i_{N+1}r}{\bm{e}}^{(N+1)}_{i_{N+1}}\right),$
which can be expressed in matrix form
$m\!\bigl{(}{\bm{x}}^{(1)},\dots,{\bm{x}}^{(N)}\bigr{)}\;\approx\;{\bm{A}}\left({\bm{B}}^{(1)\top}{\bm{x}}^{(1)}\odot\cdots\odot{\bm{B}}^{(N)\top}{\bm{x}}^{(N)}\right)$
where $\odot$ denotes the Hadamard product.
Appendix B CP-Based Tensor Product
B.1 Error Bound of CP-Based Tensor Product
{theorem}
Equivariance Error Bound of CP Decompositionequivariance-error Let CG tensor ${\bm{M}}\in\mathbb{R}^{d\times d\times d}$ and $\widehat{{\bm{M}}}$ be the rank-$R$ CP-decomposition-based approximation obtained by Frobenius minimization. For any rotation ${\bm{R}}\in SO(3)$ and any bounded $\rm{SO}(3)$-irreps ${\bm{x}},{\bm{y}}\in\mathbb{R}^{d},\lVert{\bm{x}}\rVert,\lVert{\bm{y}}\rVert\leq C$, we have
[…]
where $R_{T}=\lceil R^{1/3}\rceil$ and $\sigma_{k}^{(n)}$ is the $k$-th singular value of mode-$n$ matricization of ${\bm{M}}$.
Proof.
Given any rotation ${\bm{R}}$ and $\rm{SO}(3)$-irreps ${\bm{x}},{\bm{y}}$,
$\lVert D({\bm{R}}){\bm{x}}\rVert=\sqrt{{\bm{x}}^{\top}D({\bm{R}})^{*}D({\bm{R}}){\bm{x}}}=\sqrt{{\bm{x}}^{\top}{\bm{I}}{\bm{x}}}=\sqrt{{\bm{x}}^{\top}{\bm{x}}}=\lVert{\bm{x}}\rVert,$
and
$\begin{split}\lVert(D({\bm{R}})\otimes D({\bm{R}}))({\bm{x}}\otimes{\bm{y}})\rVert_{F}&=\sqrt{\operatorname{Tr}(({\bm{x}}\otimes{\bm{y}})^{\top}(D({\bm{R}})\otimes D({\bm{R}}))^{*}(D({\bm{R}})\otimes D({\bm{R}}))({\bm{x}}\otimes{\bm{y}}))}\\ &=\sqrt{\operatorname{Tr}(({\bm{x}}\otimes{\bm{y}})^{\top}(D({\bm{R}})^{*}D({\bm{R}})\otimes D({\bm{R}})^{*}D({\bm{R}}))({\bm{x}}\otimes{\bm{y}}))}\\ &=\sqrt{\operatorname{Tr}(({\bm{x}}\otimes{\bm{y}})^{\top}({\bm{I}}\otimes{\bm{I}})({\bm{x}}\otimes{\bm{y}}))}\\ &=\sqrt{\operatorname{Tr}(({\bm{x}}\otimes{\bm{y}})^{\top}({\bm{x}}\otimes{\bm{y}}))}\\ &=\lVert{\bm{x}}\otimes{\bm{y}}\rVert_{F}\end{split}$ (5)
Since the tensor product Frobenius norm is equal to multiplication of Frobenius norms, we have
$\begin{split}\varepsilon({\bm{R}},{\bm{x}},{\bm{y}})&=\lVert\widehat{{\bm{M}}}(D({\bm{R}}){\bm{x}}\otimes D({\bm{R}}){\bm{y}})-D({\bm{R}})\widehat{{\bm{M}}}({\bm{x}}\otimes{\bm{y}})\rVert\\ &=\lVert\widehat{{\bm{M}}}(D({\bm{R}}){\bm{x}}\otimes D({\bm{R}}){\bm{y}})-{\bm{M}}(D({\bm{R}}){\bm{x}}\otimes D({\bm{R}}){\bm{y}})\\ &\quad+D({\bm{R}}){\bm{M}}({\bm{x}}\otimes{\bm{y}})-D({\bm{R}})\widehat{{\bm{M}}}({\bm{x}}\otimes{\bm{y}})\rVert\\ &\leq\lVert\widehat{{\bm{M}}}(D({\bm{R}}){\bm{x}}\otimes D({\bm{R}}){\bm{y}})-{\bm{M}}(D({\bm{R}}){\bm{x}}\otimes D({\bm{R}}){\bm{y}})\rVert\\ &\quad+\lVert D({\bm{R}}){\bm{M}}({\bm{x}}\otimes{\bm{y}})-D({\bm{R}})\widehat{{\bm{M}}}({\bm{x}}\otimes{\bm{y}})\rVert\\ &=\lVert(\widehat{{\bm{M}}}-{\bm{M}})(D({\bm{R}}){\bm{x}}\otimes D({\bm{R}}){\bm{y}})\rVert\\ &\quad+\lVert D({\bm{R}})({\bm{M}}-\widehat{{\bm{M}}})({\bm{x}}\otimes{\bm{y}})\rVert\\ &=\lVert(\widehat{{\bm{M}}}-{\bm{M}})(D({\bm{R}})\otimes D({\bm{R}}))({\bm{x}}\otimes{\bm{y}})\rVert\\ &\quad+\lVert D({\bm{R}})({\bm{M}}-\widehat{{\bm{M}}})({\bm{x}}\otimes{\bm{y}})\rVert\\ &\leq\lVert\widehat{{\bm{M}}}-{\bm{M}}\rVert_{F}\lVert(D({\bm{R}})\otimes D({\bm{R}}))({\bm{x}}\otimes{\bm{y}})\rVert_{F}\\ &\quad+\lVert{\bm{M}}-\widehat{{\bm{M}}}\rVert_{F}\lVert{\bm{x}}\otimes{\bm{y}}\rVert_{F}\\ &=2\lVert\widehat{{\bm{M}}}-{\bm{M}}\rVert_{F}\lVert{\bm{x}}\otimes{\bm{y}}\rVert_{F}\\ &=2\lVert\widehat{{\bm{M}}}-{\bm{M}}\rVert_{F}\lVert{\bm{x}}\rVert\lVert{\bm{y}}\rVert\\ &\leq 2C^{2}\Bigl{(}\sum_{n=1}^{3}\sum_{k>R_{T}}\sigma_{k}^{2}\Bigr{)}^{1/2}\end{split}$ (6)
B.2 Universality of CP-Based Tensor Product
{theorem}
Universality of CP Decompositionuniversality Consider $\rm{SO}(3)$-irreps ${\bm{x}}\in V_{1}=\mathbb{R}^{d}$, ${\bm{y}}\in V_{2}=\mathbb{R}^{d_{2}}$ and $V_{3}=\mathbb{R}^{d_{3}}$. Fo any $\rm{SO}(3)$-equivariant bilinear map $\Phi$, there exist ${\bm{B}}\in\mathbb{R}^{d_{1}\times R}$, ${\bm{C}}\in\mathbb{R}^{d_{2}\times R}$, ${\bm{A}}\in\mathbb{R}^{R\times d_{3}}$ such that $\Phi$ can be written as
$\Phi({\bm{x}},{\bm{y}})={\bm{A}}({\bm{B}}^{\top}{\bm{x}}\odot{\bm{C}}^{\top}{\bm{y}})\in V_{3},$
if $R$ is sufficient large.
Proof.
A bilinear map $\Phi$ is uniquely encoded by a third-order tensor ${\bm{T}}\in V_{3}\otimes V_{1}\otimes V_{2}$ via $\Phi({\bm{x}},{\bm{y}})={\bm{T}}({\bm{x}}\otimes{\bm{y}})$ and all equivariant tensors ${\bm{T}}$ form the fixed-point subspace $\mathcal{H}=(V_{3}\otimes V_{1}\otimes V_{2})^{\rm{SO}(3)}=\operatorname{Hom}_{\rm{SO}(3)}(V_{1}\otimes V_{2},V_{3})$. Write each irrep $V_{i}$ as a direct sum of irreps
$V_{i}\cong\bigoplus_{\ell=0}^{L}c_{i}^{(\ell)}H^{(\ell)}$
where $c_{i}^{(\ell)}$ is the multiplicity of the irrep of degree $\ell$ in $V_{i}$ and $H^{(\ell)}$ is the $(2\ell+1)$-dimensional irrep. For every admissible path $(\ell_{1},\ell_{2},\ell_{3})$ and multiplicities $1\leq v\leq c_{1}^{(\ell_{1})},1\leq w\leq c_{2}^{(\ell_{2})},1\leq u\leq c_{3}^{(\ell_{3})}$, we pick orthonormal basis vectors ${\bm{a}}_{u}^{(\ell_{3})}\in V_{3}$, ${\bm{b}}_{v}^{(\ell_{1})}\in V_{1}$, ${\bm{c}}_{w}^{(\ell_{2})}\in V_{2}$ and define
${\bm{T}}^{(\ell_{1},\ell_{2},\ell_{3})}_{uvw}={\bm{a}}_{u}^{(\ell_{3})}\otimes{\bm{b}}_{v}^{(\ell_{1})}\otimes{\bm{c}}_{w}^{(\ell_{2})}$
so that $\{{\bm{T}}^{(\ell_{1},\ell_{2},\ell_{3})}_{uvw}\}$ is a basis of $\mathcal{H}$. And then we zero-pad ${\bm{a}}^{(\ell_{3})}_{u},{\bm{b}}^{(\ell_{1})}_{v},{\bm{c}}^{(\ell_{2})}_{w}$ as the vectors in $V_{3}$, $V_{1}$, and $V_{2}$, respectively. Now let
$\mu=\operatorname{dim}\mathcal{H}=\sum_{\ell_{1}=0}^{L}\sum_{\ell_{2}=0}^{L}\sum_{\ell_{3}=|\ell_{1}-\ell_{2}|}^{\ell_{1}+\ell_{2}\leq L}c_{1}^{(\ell_{1})}c_{2}^{(\ell_{2})}c_{3}^{(\ell_{3})},$
and rename the basis as $\{{\bm{T}}^{(k)}\}_{k=1}^{\mu}$ and vectors ${\bm{a}}^{(\ell_{3})}_{u},{\bm{b}}^{(\ell_{1})}_{v},{\bm{c}}^{(\ell_{2})}_{w}$ as ${\bm{a}}_{k},{\bm{b}}_{k},{\bm{c}}_{k}$ for the corresponding ${\bm{T}}^{(k)}={\bm{T}}^{(\ell_{1},\ell_{2},\ell_{3})}_{uvw}$. Given an arbitrary equivariant tensor ${\bm{T}}\in\mathcal{H}$, we expand it in that basis ${\bm{T}}=\sum_{k=1}^{\mu}\lambda_{k}{\bm{T}}^{(k)}$ and stack the vectors correspondingly as
${\bm{A}}=\bigl{[}\lambda_{1}{\bm{a}}_{1},\dots,\lambda_{\mu}{\bm{a}}_{\mu}\bigr{]}^{\!\top},\quad{\bm{B}}=\bigl{[}{\bm{b}}_{1},\dots,{\bm{b}}_{\mu}\bigr{]},\quad{\bm{C}}=\bigl{[}{\bm{c}}_{1},\dots,{\bm{c}}_{\mu}\bigr{]}.$
For all ${\bm{x}},{\bm{y}}$ one then has $\Phi({\bm{x}},{\bm{y}})={\bm{A}}({\bm{B}}^{\!\top}{\bm{x}}\odot{\bm{C}}^{\!\top}{\bm{y}}).$ Taking $R\geq\mu$ supplies enough columns/rows to hold this equation; any surplus columns can be zero-padded. ∎
Appendix C Model and Training Configurations
PubChemQCR training.
All experiments on the PubChemQCR benchmarks use a cutoff radius of 4.54.54.5\,4.5Å, the Adam optimizer with initial learning rate $5\times 10^{-4}$, and a ReduceLROnPlateau scheduler with a patience of 2 epochs. Models are trained for up to 100 epochs on PubChemQCR-S and up to 15 epochs on the full PubChemQCR dataset on NVIDIA A100-80GB GPUs.
TDN configuration.
For both PubChemQCR and PubChemQCR-S datasets, TDN employs 6 graph-transformer layers with MLP attention, an irreps-channel dimension of 256, a maximum angular degree $L=1$, and a graph-transformer layer for the force output head. Also, the model is trained with batch size 32 under the above training protocol. For the OC20 IS2RE-Direct task, TDN adopts the same model configuration. For the OC20 S2EF task, TDN uses a similar configuration but with 12 graph-transformer layers and an additional degree-$1$ bond-order orientation (BOO) feature [41] adding to the initial node embedding. For the OC22 IS2RE task, TDN also uses a similar configuration but with 6 graph-transformer layers and an additional degree-$9$ BOO feature [41] adding to the initial node embedding.
Table 6: Configurations including layer counts, hidden (maximum irrep-channel) dimensions, and batch sizes of baseline models including SchNet [12], PaiNN [4], MACE [7], Equiformer [6], PACE [29], FAENet [8], NequIP [20], and Allegro [31] for PubChemQCR experiments.
Model Layers Hidden Dimension Batch Size
SchNet 4 128 128
PaiNN 4 128 32
FAENet 4 128 64
NequIP 5 64 16
SevenNet 5 128 16
MACE 2 128 8
PACE 2 128 8
Allegro 2 128 8
Equiformer 4 128 32
Other baselines.
Table 6 summarizes the configurations for all other baseline models. For FAENet we select the “simple” message-passing variant; for MACE we include the residual interaction block. Initial attempts to train Equiformer with its OC20 original settings (6 layers, hidden irreps 256x0e+256x1e or 256x0e+128x1e) failed to converge; we therefore adopt a reduced configuration of 4 layers with irreps 128x0e+64x1e. For all tensor-product-based methods including NequIP [20], MACE [7], PACE [29], Allegro [31], SevenNet [30], and Equiformer [6], only even-parity irreps are retained and $L_{max}=2$ except for Equiformer.
Appendix D Limitations
Our approach accelerates $\rm{SO}(3)$-equivariant tensor products by replacing the full CG tensor product with a low-rank CP decomposition. We substantiate both its theoretical speed-up and its empirical efficacy across multiple benchmarks. A principal limitation is that the minimal rank required to achieve a desired approximation error is not known and determining it exactly is NP-hard. To address this for the CG product, we introduce an empirical rank scheduler; however, this scheduler relies on the CG tensor and must be re-derived for each new equivariant tensor product. In future work, we will investigate the general relationship between group-equivariant tensor products and their optimal rank profiles, with the goal of developing broadly applicable rank-selection strategies.